Allele for dark spots is predominant over the allele for earthy colored spots in cockroaches

They are in Hardy-Weinberg equilibrium for their locus.

98/111 had dark spots

13/111 had earthly colored spots.

1. How many cockroaches in this group are most expected to be heterozygous at the locus?

2. Give the frequency of the allele for the earthly colored spots.

3. Give the frequency of the allele for dark spots.
Answer #1

Given,98/111 have dark spots and 13/111 have earthy spot. Let p be allele frequency of dominant trait and q be allele frequency of recessive trait.

Let calculate number of recessive individuals if population size is 100.Number of recessive individuals=(13/111)*100=11.71. So,q^2= 11.71/100=0.1171 and q= 0.342.

Then p=1-0.342=0.658.

Number of homozygous dominant individuals = p^2*100=0.658^2=0.433*100=43.3.

For population of 111,this number is [(111*43.3)/100] = 48.0 individuals.

Number of heterozygous dominant individuals = 2*p*q*100=2*0.658*0.342*100=45.0.

For population of 111,this number is[(111*45)/100]=49.95=50 individuals.

1.Number of cockroaches in this group are most expected to be heterozygous at the locus=50 individuals.

2.Allele freqency for the earthly colored spots(q) = 0.342

3.Allele frequency for dark spots(p)=0.658.